a truck is moving with a speed of 3 m/s is accelerated so that its speed achieved 8 m/s after 5 second. Then the driver hit the brake and the truck stop within 7 seconds. Determine the total distance travelled by the truck


1
Expert's answer
2020-09-28T08:06:07-0400

The total distance travelled by the truck

"d=d_1+d_2"

Here

"d_1=\frac{v_1+v_2}{2}t_1=\rm\frac{3\: m/s+5\: m/s}{2}\times 5\: s=20\: m"

"d_2=\frac{v_2+v_3}{2}t_2=\rm\frac{5\: m/s+0\: m/s}{2}\times 7\: s=17.5\: m"

Finally

"d=d_1+d_2=20\:\rm m+17.5\: m=37.5\: m"